# Long-term impact of the magnetic-field strength on the evolution and electromagnetic emission by neutron-star merger remnants

### v1

@@ I Introduction
The second aspect involves an urgent need for long-term, large-scale simulations to consistently capture all multi-messenger signals produced in compact binary merger events, e.g., GWs Abbott et al. (2017, 2019), GRB multi-band EM emission Fermi-LAT Collaboration (2017); Goldstein et al. (2017), multi-band afterglow emission Hallinan et al. (2017); Troja et al. (2017); Margutti et al. (2018); Mooley et al. (2018); Ghirlanda et al. (2019); Musolino et al. (2024b), and kilonova observations Chornock and et al. (2017); Li and Paczyński (1998); Drout et al. (2017); Nicholl et al. (2017); Pian et al. (2017); Soares-Santos et al. (2017). These general-relativistic simulations must cover timescales of seconds to account for the delay from the GW to the associated GRB signal, which is $1.7\,\mathrm{s}$ in the case of GW170817 Goldstein et al. (2017), and hints to a collapse time of about $1.0\,\mathrm{s}$ Gill et al. (2019); Murguia-Berthier et al. (2021); Nathanail et al. (2021a). Even longer timescales, although not necessarily in full general relativity (GR), must be span to describe the afterglow signal appearing days later Margutti et al. (2017) and lasting months Troja et al. (2019). This presents challenges in maintaining the same simulation method over extended periods. Hence, transitions between codes (hand-off) at different stages offer a versatile approach, enabling various physical processes to be modelled while leveraging the unique strengths of different codes (see, e.g.,  Lopez Armengol et al. (2022); Gottlieb et al. (2023a, b, 2023); Ng et al. (2024a); Ennoggi et al. (2025)).
@@
Using the Conformal Flatness Condition (CFC) to solve the elliptic sector of the Einstein equations offers an efficient approach to these challenges, as it reduces the need for frequent spacetime computations when the spacetime is not highly dynamic Ng et al. (2024a). The CFC approximation has been successfully applied in simulations of core-collapse supernovae Dimmelmeier et al. (2002a); Ott et al. (2007); Müller (2015); Cheong et al. (2023); Ng et al. (2024b), rapidly rotating neutron stars Dimmelmeier et al. (2002b); Cordero-Carrión et al. (2009); Ng et al. (2021); Cheong et al. (2024a), and BNS mergers Bauswein et al. (2012, 2021); Lioutas et al. (2024). Our previous work Ng et al. (2024a), further demonstrated that the extended-CFC (xCFC) scheme Cordero-Carrión et al. (2009) implemented in our new , multi-coordinate and multi-dimensional GRMHD code BHAC+ can effectively handle the output of another code, FIL Most et al. (2019, 2021); Chabanov et al. (2023); Musolino and Rezzolla (2024); Chabanov and Rezzolla (2023); Ng et al. (2024), which solves the Einstein equations using full general relativity. Specifically, we have shown that the CFC approximation provides nearly identical solutions in the head-on collision of two neutron stars, not only in the evolution of maximum rest-mass density, but also in the two-dimensional (2D) rest-mass density and temperature distributions, when compared to the full-GR code FIL.
@@ II.2 GW radiation-reaction in the xCFC scheme
Some of the initial work on the inclusion of RR on the energy and momentum budget in numerical codes solving the post-Newtonian (PN) formulation of the hydrodynamic equations was reported in Refs. Blanchet et al. (1990); Rezzolla et al. (1999); Faye and Schäfer (2003). These approaches where then coupled phenomenologically with the CFC scheme thereby compensating for the waveless property of the CFC approximation Oechslin et al. (2007); Lioutas et al. (2024). By implementing the 3.5 PN order formalism, but including only all 2.5PN terms and 3.5 PN corrections related to the gravitational potential and its derivatives in the metric components, the RR scheme provides a gravitational description with modified metric components coupled to hydrodynamics to approximate the system’s energy-momentum loss via GWs.
@@ III.1 FIL, FUKA and BHAC+
FIL Most et al. (2019, 2021); Chabanov et al. (2023); Musolino and Rezzolla (2024); Chabanov and Rezzolla (2023); Ng et al. (2024) solves the hyperbolic sector of the Einstein equations Alcubierre (2006); Rezzolla and Zanotti (2013) using full GR schemes, including BSSNOK Shibata and Nakamura (1995); Baumgarte and Shapiro (1999), BSSNOK-Z4c Bernuzzi and Hilditch (2010), and CCZ4 Alic et al. (2012, 2013). Coupled with the EinsteinToolkit Loeffler et al. (2012); Zlochower and et al. (2022), it utilises the Carpet box-in-box Adaptive Mesh Refinement (AMR) driver in Cartesian coordinates Schnetter et al. (2006). Besides, it uses a fourth-order accurate conservative finite-difference High-Resolution Shock-Capturing (HRSC) scheme with WENO-Z reconstruction Del Zanna et al. (2007), an HLL Riemann solver, and a vector-potential-based method Etienne et al. (2015).
@@
In the default setup, BHAC+ makes use of the WENO-Z reconstruction and fourth-order finite differencing and this leads to the closest matches with the results from FIL in terms of the evolution of maximum density, rotational profiles, and total EM energy. It is important to mention that both approaches to the violation of the divergence-free condition, namely, the use of a vector potential or of the constrained-transport method Evans1988, guarantee a magnetic field that is divergence free at machine precision. It is also worth noting that we have resolved an artefact reported in our previous work Ng et al. (2024a), where we observed that when performing an azimuthally averaged hand-off, the postmerger remnant experienced fluctuating and significantly lower temperatures in the core region. As discussed in Ng et al. (2024a), this behaviour was partly due to slightly inaccurate values of the specific internal energy $\epsilon$ generated by metric initialisation with the azimuthally averaged hand-off, but also to the poor resolution of the tabulated EOS, which has a strong dependence of the temperature $T$ on the specific internal energy $\epsilon$ in these high-density regimes. Although these artefacts are reduced when employing a full 3D hand-off, a series of experiments has revealed that reconstructing the $T$ instead of the $\epsilon$ during evolution leads to significant improvements of this issue and is effective across various EOSs.
@@ III.2.2 High-order interpolation for spacetime and matter quantities
Tests baryonic mass EM energy int. energy
@@ IV Validation of the hand-off and RR
The CFC approach has been successfully applied in the study of uniformly and differentially rotating neutron stars Cook et al. (1996); Iosif and Stergioulas (2014, 2021); Cheong et al. (2024a, b), with only minor deviations observed compared to the full-GR solutions. Furthermore, our previous work Ng et al. (2024a) has demonstrated that the CFC approach can closely approximate the full-GR solution in the highly nonlinear system of neutron stars head-on collisions in full 3D simulations. However, the performance of the CFC in realistic BNS postmerger scenarios with magnetic fields and its comparison with full-GR simulations, has not been explored thus far.
@@ V.1 Impact on the EM-energy evolution
Both the toroidal and poloidal EM energy grow rapidly during the KHI stage, peaking at around $\bar{t}\approx 2\,\rm ms$. After this point, the EM energy begins to rapidly decrease, mostly due to the lack of sufficient turbulent vortices to counteract dissipation (see discussion in Chabanov et al. (2023)). As the merger remnant develops a large-scale differential rotation structure, the winding stage commences and the toroidal magnetic-field strength exhibits the characteristic linear growth over time, which can last up to $\simeq 30\,{\rm ms}$ ($\simeq 90\,{\rm ms}$) for the HMF (LMF) case at high resolution. Over the same period of time, we observe a rapid decay in the poloidal-energy component, primarily due to numerical dissipation caused by a combination of shocks and violent oscillations within the remnant.
Magnetic winding ceases when the toroidal magnetic fields become strong enough to exert an effective torque on the fluid, resulting in angular momentum redistribution. This transition, which we denominate as “slow-down” occurs earlier in the HMF case (i.e., $\simeq 30\,{\rm ms}$) than in the LMF case ($\simeq 90\,{\rm ms}$); the relevant times are marked with vertical dashed lines in Fig. 5. It is important to underline that the slow-down does not mark the end of the magnetic-field winding, which indeed continues also afterwards, but rather the beginning of a different regime in which winding of the poloidal magnetic field into the toroidal one still takes place but is slowed down, i.e., it leads to a sublinear in time growth, because of the intense back-reaction of the magnetic tension (see also the discussion in Sect. V.2). In addition, we should note that we expect the specific threshold depends, in addition to the initial magnetic-field strength, also on the rotational profile of the merger remnants and hence also on the mass ratio and EOS. This redistribution of angular momentum simultaneously generates an effective poloidal fluid motion, counteracting the dissipation of poloidal EM energy. Additionally, there is another contribution to the poloidal EM-energy growth after the EM amplification ceases and that comes from buoyant toroidal field lines that rise to higher latitudes, causing magnetic flares that build structured poloidal magnetic fields in the polar region (see discussion below).
@@ V.3 Parker instability and flares
By the end of all simulations, and both in the LMF and HMF scenarios, the merger remnant becomes approximately spherical (see density contours in Fig. 8), partly due to the angular-momentum loss and partly due to the reduced pinch-effect of the strong toroidal magnetic field that produces a prolate distribution of matter Frieben and Rezzolla (2012). The reduced deformation of the magnetar is also responsible for a reduction in the amplitude of the remnant oscillations and for the dynamical ejection of matter, both in the polar and equatorial directions. As a result, matter will tend to accumulate and, in the absence of a neutrino wind sweeping away to large distances, it will lead to a change in the stratification in the outer layers of the HMNS and hence to a reduced violation of the Parker instability criterion. In turn, this results into a partial weakening of the Poynting flux around $\sim 200\,\rm ms$ (see Figs. 7 and 11, and the discussion below).
@@ V.4 Properties of the outflow
Also reported in Fig. 11, but in the lower panel, is the evolution of the ejected mass (see Appendix A for details on the definition and calculation). Clearly, a substantial amount of mass is ejected by the remnant over the timescale of the simulation and this amounts to a total of $M_{\rm ej}\sim 0.11\,M_{\odot}$ for the LMF case and $\sim 0.17\,M_{\odot}$ for the HMF case, thus indicating that strong magnetic fields in the merger remnant greatly facilitate the mass ejection. Less obvious is why the for $20\lesssim\bar{t}\lesssim 50\,{\rm ms}$ the ejected mass is larger in the LMF case than in the HMF case. We believe that this is because of the larger braking action in the HMF case, that converts the kinetic-energy reservoir in the remnant into toroidal magnetic field. This braking is smaller for the LMF scenario and hence the kinetic energy is employed more efficiently to eject unbound matter. Part of the reason may stem from the pinch effect of the strong toroidal magnetic field (see, e.g.,  Frieben and Rezzolla (2012)), which has already reached very large values by $\bar{t}\sim 20\,\rm{ms}$ in the HMF case. This effect can limit matter expansion in the disk region, the primary source of the ejected mass.
@@ V.5 Dynamics of the flares near the HMNS
At small radii, i.e., $r\sim 15\,\rm{km}$, the strong Poynting flux primarily originates from flares, namely, matter outflows carrying intense magnetic fields during the violent oscillation phase of the HMNS. At the eruption, magnetic energy in the flare is converted to kinetic energy but, as they propagate outward, they naturally interact with the ambient plasma, transferring part of their kinetic energy, but also expanding and cooling, and ultimately weakening their strength, as a result, the Poynting flux at large radii, i.e., $r\sim 300\,\rm{km}$, has been reduced significantly. Furthermore, we observe distinct morphological differences across the different two-spheres: in the polar regions, structures tend to form ring-like patterns, whereas at lower latitudes, they resemble belts. This distinction arises naturally due to the rapid rotation of the remnant and the large degree of anisotropy in the ejection of matter.
@@ VI Conclusions
More specifically, using these improvements, we have stress-tested the hybrid approach carrying out the hand-off procedure under different conditions of the spacetime dynamics and GW content. Our findings indicate that this approach yields highly accurate evolutions for the postmerger remnant, capturing both overall dynamics and magnetic field evolution effectively. The agreement between the two codes greatly improves as the GW amplitude decreases, especially when GW amplitude has decreased to $1/16$ of its peak value.
@@ Appendix E On the conformal flatness of BNS postmerger spacetimes
Furthermore, and as expected, during the merger all three measures increase rapidly. The quantity $\langle H\rangle_{\rho}$ achieves moderate values of $\sim 10^{-1}$, which agree with strongly differentially rotating relativistic stars Iosif and Stergioulas (2014). More specifically, these values correspond broadly to the configurations with the highest values of $T/|W|$ in Ref. Iosif and Stergioulas (2014), where $T/|W|$ denotes the ratio of the rotational over the gravitational binding energy; with such values of $T/|W|$, we expect the remnant to still be stable with respect to the dynamical $\ell=2,m=2$ barmode instability Shibata and Sekiguchi (2005); Baiotti et al. (2007); Manca et al. (2007), even in the presence of strong magnetic fields Franci et al. (2013). Between merger and the time PM-A, all three measures display strong and rapid variations as a result of the violent collision-and-bounce cycles of the two stellar cores. After PM-A, the quantity $\left\langle H\right\rangle_{\rho}$ remains almost constant as a result of weak GW emission, the only mechanism to change $T/|W|$.

### v2

@@ I Introduction
The second aspect involves an urgent need for long-term, large-scale simulations to consistently capture all multi-messenger signals produced in compact binary merger events, e.g., GWs Abbott et al. (2017, 2019), GRB multi-band EM emission Fermi-LAT Collaboration (2017); Goldstein et al. (2017), multi-band afterglow emission Hallinan et al. (2017); Troja et al. (2017); Margutti et al. (2018); Mooley et al. (2018); Ghirlanda et al. (2019); Musolino et al. (2024b), and kilonova observations Chornock and et al. (2017); Li and Paczyński (1998); Drout et al. (2017); Nicholl et al. (2017); Pian et al. (2017); Soares-Santos et al. (2017). These general-relativistic simulations must cover timescales of seconds to account for the delay from the GW to the associated GRB signal, which is $1.7\,\mathrm{s}$ in the case of GW170817 Goldstein et al. (2017), and hints at a collapse time of about $1.0\,\mathrm{s}$ Gill et al. (2019); Murguia-Berthier et al. (2021); Nathanail et al. (2021a). Even longer timescales, although not necessarily in full general relativity (GR), must be span to describe the afterglow signal appearing days later Margutti et al. (2017) and lasting months Troja et al. (2019). This presents challenges in maintaining the same simulation method over extended periods. Hence, transitions between codes (hand-off) at different stages offer a versatile approach, enabling various physical processes to be modelled while leveraging the unique strengths of different codes (see, e.g.,  Lopez Armengol et al. (2022); Gottlieb et al. (2023a, b, 2023); Ng et al. (2024a); Ennoggi et al. (2025)).
@@
Using the Conformal Flatness Condition (CFC) to solve the elliptic sector of the Einstein equations offers an efficient approach to these challenges, as it reduces the need for frequent spacetime computations when the spacetime is not highly dynamic Ng et al. (2024a). The CFC approximation has been successfully applied in simulations of core-collapse supernovae Dimmelmeier et al. (2002a); Ott et al. (2007); Müller (2015); Cheong et al. (2023); Ng et al. (2024b), rapidly rotating neutron stars Dimmelmeier et al. (2002b); Cordero-Carrión et al. (2009); Ng et al. (2021); Cheong et al. (2024), and BNS mergers Bauswein et al. (2012, 2021); Lioutas et al. (2024). Our previous work Ng et al. (2024a), further demonstrated that the extended-CFC (xCFC) scheme Cordero-Carrión et al. (2009) implemented in our new , multi-coordinate and multi-dimensional GRMHD code BHAC+ can effectively handle the output of another code, FIL Most et al. (2019, 2021); Chabanov et al. (2023); Musolino and Rezzolla (2024); Chabanov and Rezzolla (2023); Ng et al. (2024), which solves the Einstein equations using full general relativity. Specifically, we have shown that the CFC approximation provides nearly identical solutions in the head-on collision of two neutron stars, not only in the evolution of maximum rest-mass density, but also in the two-dimensional (2D) rest-mass density and temperature distributions, when compared to the full-GR code FIL.
@@ II.2 GW radiation-reaction in the xCFC scheme
Some of the initial work on the inclusion of RR on the energy and momentum budget in numerical codes solving the post-Newtonian (PN) formulation of the hydrodynamic equations was reported in Refs. Blanchet et al. (1990); Rezzolla et al. (1999); Faye and Schäfer (2003). These approaches were then coupled phenomenologically with the CFC scheme thereby compensating for the waveless property of the CFC approximation Oechslin et al. (2007); Lioutas et al. (2024). By implementing the 3.5 PN order formalism, but including only all 2.5PN terms and 3.5 PN corrections related to the gravitational potential and its derivatives in the metric components, the RR scheme provides a gravitational description with modified metric components coupled to hydrodynamics to approximate the system’s energy-momentum loss via GWs.
@@ III.1 FIL, FUKA and BHAC+
FIL Most et al. (2019, 2021); Chabanov et al. (2023); Musolino and Rezzolla (2024); Chabanov and Rezzolla (2023); Ng et al. (2024) solves the hyperbolic sector of the Einstein equations Alcubierre (2006); Rezzolla and Zanotti (2013) using full GR schemes, including BSSNOK Shibata and Nakamura (1995); Baumgarte and Shapiro (1999), BSSNOK-Z4c Bernuzzi and Hilditch (2010), and CCZ4 Alic et al. (2012, 2013). Coupled with the EinsteinToolkit Loeffler et al. (2012); Zlochower and et al. (2022), it utilises the Carpet box-in-box Adaptive Mesh Refinement (AMR) driver in Cartesian coordinates Schnetter et al. (2006). Besides, it uses a fourth-order accurate conservative finite-difference High-Resolution Shock-Capturing (HRSC) scheme with WENO-Z reconstruction Del Zanna et al. (2007), an HLL Riemann solver, and a vector-potential-based magnetic evolving method Etienne et al. (2015).
@@
In the default setup, BHAC+ makes use of the WENO-Z reconstruction and fourth-order finite differencing and this leads to the closest matches with the results from FIL in terms of the evolution of maximum density, rotational profiles, and total EM energy. It is important to mention that both approaches to the violation of the divergence-free condition, namely, the use of a vector potential or of the constrained-transport method Evans and Hawley (1988), guarantee a magnetic field that is divergence free at machine precision. It is also worth noting that we have resolved an artefact reported in our previous work Ng et al. (2024a), where we observed that when performing an azimuthally averaged hand-off, the postmerger remnant experienced fluctuating and significantly lower temperatures in the core region. As discussed in Ng et al. (2024a), this behaviour was partly due to slightly inaccurate values of the specific internal energy $\epsilon$ generated by metric initialisation with the azimuthally averaged hand-off, but also to the poor resolution of the tabulated EOS, which has a strong dependence of the temperature $T$ on the specific internal energy $\epsilon$ in these high-density regimes. Although these artefacts are reduced when employing a full 3D hand-off, a series of experiments has revealed that reconstructing the $T$ instead of the $\epsilon$ during evolution leads to significant improvements of this issue and is effective across various EOSs.
@@ III.2.2 High-order interpolation for spacetime and matter quantities
Tests Baryonic mass EM energy Internal energy
@@ IV Validation of the hand-off and RR
The CFC approach has been successfully applied in the study of uniformly and differentially rotating neutron stars Cook et al. (1996); Iosif and Stergioulas (2014, 2021); Cheong et al. (2024, 2025), with only minor deviations observed compared to the full-GR solutions. Furthermore, our previous work Ng et al. (2024a) has demonstrated that the CFC approach can closely approximate the full-GR solution in the highly nonlinear system of neutron stars head-on collisions in full 3D simulations. However, the performance of the CFC in realistic BNS postmerger scenarios with magnetic fields and its comparison with full-GR simulations, has not been explored thus far.
@@ IV.1 On the role of hand-off time
In summary, our results show that using the GW amplitude represents a gauge-invariant, model-independent, and essentially resolution-independent manner to assess when the hand-off can be performed. In addition, we suggest that performing the hand-off when the GW amplitude has reached $\simeq 1/16$ of the peak amplitude represents a robust choice. While these conclusions are based on a limited number of tests, the underlying motivation is clear: the smaller the GW emission, the better the CFC approximation and its use at hand-off. At the same time, we also recommend that some experimentation is made around the hand-off time. While this experimentation brings in some extra computational costs, the latter are minute when compared with those associated with the long-term evolution that can be performed under optimal hand-off conditions.
@@ V.1 Impact on the EM-energy evolution
Both the toroidal and poloidal EM energy grow rapidly during the KHI stage, peaking at around $\bar{t}\approx 2\,\rm ms$. After this point, the EM energy begins to rapidly decrease, mostly due to the lack of sufficient turbulent vortices to counteract dissipation (see discussion in Chabanov et al. (2023)). As the merger remnant develops a large-scale differential rotation structure, the winding stage commences and the toroidal magnetic-field strength exhibits the characteristic linear growth over time, which can last up to $\bar{t}\simeq 30\,{\rm ms}$ ($\bar{t}\simeq 90\,{\rm ms}$) for the HMF (LMF) case at high resolution. Over the same period of time, we observe a rapid decay in the poloidal-energy component, primarily due to numerical dissipation caused by a combination of shocks and violent oscillations within the remnant.
Magnetic winding ceases when the toroidal magnetic fields become strong enough to exert an effective torque on the fluid, resulting in angular momentum redistribution. This transition, which we denominate as “slow-down” occurs earlier in the HMF case (i.e., $\bar{t}\simeq 30\,{\rm ms}$) than in the LMF case ($\bar{t}\simeq 90\,{\rm ms}$); the relevant times are marked with vertical dashed lines in Fig. 5. It is important to underline that the slow-down does not mark the end of the magnetic-field winding, which indeed continues also afterwards, but rather the beginning of a different regime in which winding of the poloidal magnetic field into the toroidal one still takes place but is slowed down, i.e., it leads to a sublinear in time growth, because of the intense back-reaction of the magnetic tension (see also the discussion in Sect. V.2). In addition, we should note that we expect the specific threshold depends, in addition to the initial magnetic-field strength, also on the rotational profile of the merger remnants and hence also on the mass ratio and EOS. This redistribution of angular momentum simultaneously generates an effective poloidal fluid motion, counteracting the dissipation of poloidal EM energy. Additionally, there is another contribution to the poloidal EM-energy growth after the EM amplification ceases and it comes from buoyant toroidal field lines that rise to higher latitudes, causing magnetic flares that build structured poloidal magnetic fields in the polar region (see discussion below).
@@ V.3 Parker instability and flares
By the end of all simulations, and both in the LMF and HMF scenarios, the merger remnant becomes approximately spherical (see density contours in Fig. 8), partly due to the angular-momentum loss and partly due to the reduced pinch-effect of the strong toroidal magnetic field that produces a prolate distribution of matter Frieben and Rezzolla (2012). The reduced deformation of the magnetar is also responsible for a reduction in the amplitude of the remnant oscillations and for the dynamical ejection of matter, both in the polar and equatorial directions. As a result, matter will tend to accumulate, and in the absence of a neutrino wind sweeping away to large distances, it will lead to a change in the stratification in the outer layers of the HMNS and hence to a reduced violation of the Parker instability criterion. In turn, this results into a partial weakening of the Poynting flux around $\sim 200\,\rm ms$ (see Figs. 7 and 11, and the discussion below).
@@ V.4 Properties of the outflow
Also reported in Fig. 11, but in the lower panel, is the evolution of the ejected mass (see Appendix A for details on the definition and calculation). Clearly, a substantial amount of mass is ejected by the remnant over the timescale of the simulation and this amounts to a total of $M_{\rm ej}\sim 0.11\,M_{\odot}$ for the LMF case and $\sim 0.17\,M_{\odot}$ for the HMF case, thus indicating that strong magnetic fields in the merger remnant greatly facilitate the mass ejection. Less obvious is why for $20\lesssim\bar{t}\lesssim 50\,{\rm ms}$ the ejected mass is larger in the LMF case than in the HMF case. We believe that this is because of the larger braking action in the HMF case, that converts the kinetic-energy reservoir in the remnant into toroidal magnetic field. This braking is smaller for the LMF scenario and hence the kinetic energy is employed more efficiently to eject unbound matter. Part of the reason may stem from the pinch effect of the strong toroidal magnetic field (see, e.g.,  Frieben and Rezzolla (2012)), which has already reached very large values by $\bar{t}\sim 20\,\rm{ms}$ in the HMF case. This effect can limit matter expansion in the disk region, the primary source of the ejected mass.
@@ V.5 Dynamics of the flares near the HMNS
At small radii, i.e., $r\sim 15\,\rm{km}$, the strong Poynting flux primarily originates from flares, namely, matter outflows carrying intense magnetic fields during the violent oscillation phase of the HMNS. At the eruption, magnetic energy in the flare is converted to kinetic energy, but as they propagate outward, they naturally interact with the ambient plasma, transferring part of their kinetic energy, but also expanding and cooling, and ultimately weakening their strength, as a result, the Poynting flux at large radii, i.e., $r\sim 300\,\rm{km}$, has been reduced significantly. Furthermore, we observe distinct morphological differences across the different two-spheres: in the polar regions, structures tend to form ring-like patterns, whereas at lower latitudes, they resemble belts. This distinction arises naturally due to the rapid rotation of the remnant and the large degree of anisotropy in the ejection of matter.
@@ VI Conclusions
More specifically, using these improvements, we have stress-tested the hybrid approach carrying out the hand-off procedure under different conditions of the spacetime dynamics and GW content. Our findings indicate that this approach yields highly accurate evolutions for the postmerger remnant, capturing both overall dynamics and magnetic field evolution effectively. The agreement between the two codes greatly improves as the GW amplitude decreases, especially when GW amplitude has decreased to $1/16$ of its peak value. We suggest the latter as the reference criterion for performing the hand-off as it is gauge-invariant, model independent, and only weakly dependent on resolution.
@@ Appendix E On the conformal flatness of BNS postmerger spacetimes
Furthermore, and as expected, during the merger all three measures increase rapidly. The quantity $\langle H\rangle_{\rho}$ achieves moderate values of $\sim 10^{-1}$, which agree with strongly differentially rotating relativistic stars Iosif and Stergioulas (2014). More specifically, these values correspond broadly to the configurations with the highest values of $T/|W|$ in Ref. Iosif and Stergioulas (2014), where $T/|W|$ denotes the ratio of the rotational over the gravitational binding energy; with such values of $T/|W|$, we expect the remnant to still be stable with respect to the dynamical $\ell=2,m=2$ barmode instability Shibata and Sekiguchi (2005); Baiotti et al. (2007); Manca et al. (2007), even in the presence of strong magnetic fields Franci et al. (2013). Between merger and the time PM-A, all three measures display strong and rapid variations as a result of the violent collision-and-bounce cycles of the two stellar cores. After PM-A, the quantity $\left\langle H\right\rangle_{\rho}$ remains almost constant as a result of weak GW emission, the dominant mechanism to change $T/|W|$.